soybean meal is 14% protein, cornmeal is 7% protein. How many pounds of each should be mixed together in order to get 280lb mixture that is 9% protein?
Answer by checkley77(12844)

.14x+.07(280-x)=.09*280
.14x+19.6-.07x=25.2
.07x=25.2-19.6
.07x=5.6
x=5.6/.07
x=80 lbs. of soybeans is used
280-80=200 lbs. of corn meal is used.
Proof:
.14*80+.07*200=25.2
11.2+14=25.2
25.2=25.2